OpenStudy (anonymous):

Find the indicated quantity, given the following vectors. u = 6i + j, v = 4i - 3j, w = 8i + 2j u · v + u · w

OpenStudy (anonymous):

24-3+48+2=71

OpenStudy (anonymous):

how do you multitply out?? Find the indicated quantity, given the following vectors. u = 7i + 5j, v = 4i - 4j, w = 8i + 4j u · (v + w)

OpenStudy (anonymous):

<7,5> <12,0> = 84, 0 ???

OpenStudy (anonymous):

so lost. :(( thanks for helping!

OpenStudy (alexwee123):